A desk top has the shape of a trapezoid. The two parallel sides have lengths 2m and 2.8 m, and the distance between the sides is 1.5m. To protect the wood surface, a blotter will be made, at a cost of $25 per square meter, to cover the entire desk top. Find the total cost of the blotter.
Answer by edjones(8007)

a=altitude, b=base, a trapezoid has 2 bases.
a((b[1]+b[2])/2)=A
=1.5((2+2.8)/2)
=1.5(4.8/2)
=1.5*2.4
=3.6 m^2
3.6*25=$90. Cost of the blotter
.
Ed